Consider the equation log(3x - 1) = log28. Explain why 3x - 1 is not equal to 8. Describe the steps you would take to solve the equation, and state what 3x - 1 is equal to
Answer by Alan3354(69443)

Consider the equation log(3x - 1) = log28. Explain why 3x - 1 is not equal to 8. Describe the steps you would take to solve the equation, and state what 3x - 1 is equal to
-------------------
log(3x-1) = log(28)
3x-1 = 28